# Minimum Cost Loop Nests for Contraction of a Sparse Tensor with a Tensor Network

### v1

@@ 5. SpTTN-Cyclops Framework
1:Input: Sequence of terms that represents the
2:          contraction path. Each term is a set
3:          of three tensors, inp1, inp2 and op.
4:Input: depth initially set to 0.
5:Output: Loop nest to compute the given tensor
6:           kernel.
7:procedure LOOP_NEST(sequence_of_terms,
8:                              depth)
@@
11:         $\text{contract}(t.\text{inp1},t.\text{inp2},t.\text{op})$
@@
20:                       $b\leftarrow\text{buf\_terms}[i]$
@@
1$\dots$
2for $q_{1}$:
3  for $q_{2}$:
4    for $q_{3}$:
5      $\dots$
6      for $q_{\alpha}$:
7        $L_{i}$
8      for $q_{\beta}$:
9        $L_{i+1}$
10    for $q_{\gamma}$:
11      $\dots$
12  for $q_{\delta}$:
13    $\dots$
14    $L_{\ast}$
15  for $q_{2}$:
16    for $q_{3}$:
17      $\dots$
18      for $q_{\varepsilon}$:
19        $L_{j-1}$
20      for $q_{\lambda}$:
21        $L_{j}$
22$\dots$
Figure 2. An illustration of the algorithm’s execution on a sequence of terms, $(L_{i},\ldots,L_{j})$. Let the indices in the sequence of terms be $(q_{1},\ldots,q_{n})$. Indices $q_{\alpha},q_{\beta},q_{\gamma},q_{\delta},q_{\varepsilon},q_{\lambda}\in(q_{1},\ldots,q_{n})$. $L_{\ast}$ is a term in the sequence nested within loops $(q_{1},q_{\delta},\ldots)$. Note that $q_{2}$ is not present in the term $L_{\ast}$. The code path (loop nest) is shown as pseudo-code without the necessary loop bounds.
1T_csf = CSF(${\mathcal{T}\_ijkl}$)
2for (i,T_i) in T_csf:
3  for (j,T_ij) in T_i:
4    $\mathcal{Y}$ = 0 // reset intermediate tensor
5    for (k,T_ijk) in T_ij:
6      $\mathcal{X}$ = 0 // reset intermediate tensor
7      for l in T_ijk:
8        xAXPY(T, t_ijkl, $\mathcal{W}$[l,:], 1, $\mathcal{X}$, 1)
9      xGER(T, S, 1, $\mathcal{X}$, 1, $\mathcal{V}$[k,:], 1, $\mathcal{Y}$, T)}
10    for r in range(R): // dense loop
11      xAXPY((S * T), $\mathcal{Y}$, $\mathcal{U}$[j,r], 1, $\mathcal{Z}$[i,r,:,:], 1)
[…]
5.1. Algorithm to Generate and Execute Loop Nests
Given a fully fused loop nest tree, in Algorithm 2 we present a runtime algorithm to generate loop nests and execute the contractions. We represent the tree with a sequence of terms (leaves) and a list per term representing the index order (vertices). This representation is sufficient for the algorithm to infer the structure of a fully fused loop nest tree.
The algorithm proceeds recursively. At each level, we group consecutive terms that have a common index at that level (subtrees) (lines 6 to 10). We generate a loop for the index and include all the terms of the subtree inside the loop (lines 22 to 23 and 27 to 28). Since the sequence of terms also represents the contraction path, the algorithm recurses over the leftmost subtree first before identifying and generating loop nests for the consecutive subtrees. In Figure 2, we show an illustration of the algorithm’s execution.
We use Algorithm 2 in two stages. In the first stage, we preprocess the fully fused loop nest tree and add hooks to (1) generate nested loops for the dense indices using metaprogramming, (2) identify independent dense loops that can be offloaded to BLAS like kernels. We also allocate memory for the intermediate tensors in this stage. In the second stage, we compute the kernel by executing the preprocessed fully fused loop nest tree. We check for hooks in Line 2 and offload the computation accordingly.
@@ 5.2. Data Distribution
5.3. End-to-end Example of SpTTN Execution
Consider the computation of an order 4 TTMc kernel,
$\mathcal{S}(i,r,s,t)=\sum_{j,k,l}\mathcal{T}(i,j,k,l)\cdot\mathcal{U}(j,r)\cdot\mathcal{V}(k,s)\cdot\mathcal{W}(l,t).$
Autotuner: The autotuner tunes over various ways to contract the tensors, $(\mathcal{T}\cdot\mathcal{U}\cdot\mathcal{V}\cdot\mathcal{W})\rightarrow\mathcal{S}$, and the corresponding indices in the contraction terms, to pick an optimal contraction path and index order. An optimal contraction path in terms of asymptotic complexity is the one that has a maximum loop depth of 5. The permutations of the index order are restricted to orders that have the sparse indices $i,j,k,l$ appear as they are stored in the CSF representation of $\mathcal{T}$. The search space is further pruned by evaluating the schedules based on the cost models presented in Section 4.2 as discussed earlier in this section.
@@ 5.3. End-to-end Example of SpTTN Execution
Data distribution: The sparse tensor is retained in the initial distribution. Tensors $\mathcal{U}$, $\mathcal{V}$, $\mathcal{W}$ and the output tensor $\mathcal{S}$ share an index each with the sparse tensor. Indices that are not shared are replicated on each processor i.e., $r\in\mathcal{U}$, $s\in\mathcal{V}$, $t\in\mathcal{W}$ and $(r,s,t)\in\mathcal{S}$ are replicated. The dimensions $I,J,K$ and $L$, which are large in size, maintain their original distribution on the processor grid.
Runtime: Finally, we use the recursive algorithm (Algorithm 2) to generate the loop nests at runtime and check for an opportunity to offload the independent dense loops based on the hooks generated in the preprocessing step. We show the final fully fused loop nest with the BLAS kernels in Figure 3. After the computation, we perform reduction on the tensor $\mathcal{S}$ and redistribute it back to its original mapping on the processor grid.
@@ 6. Related Work
In SparseLNR (Dias et al., 2022), the authors extend TACO’s scheduling space to support kernel distribution/fusion. In ReACT (Zhou et al., 2023) (implemented within the COMET compiler), a tensor expression is split into individual operations and intermediate tensors are introduced to store temporary results. Further, the outermost loops (prefix nodes) are fused resulting in a reduction of the intermediate tensor dimensions. Both SparseLNR and COMET enable the factorize-and-fuse approach. The contraction path and index orders for these loop nests are user-specified. Our main contribution is in fully enumerating the space of loop nests and finding a cost-optimal schedule automatically. Furthermore, in our evaluation (in Section 7), we show that SpTTN-Cyclops outperforms SparseLNR by orders of magnitude. For example, across various input tensors considered, SpTTN-Cyclops outperforms SparseLNR by 1.3x to 3.4x and 4x to 110.5x on MTTKRP and TTMc kernels, respectively.
Mosaic (Bansal et al., 2023) is a recent work that proposes a sparse tensor algebra compiler that can bind tensor sub-expressions to external functions of other tensor algebra libraries and compilers, and provides an automated search mechanism to find all valid bindings. In our framework, we automatically bind the innermost loops to other optimized routines (like BLAS) at runtime.
Auto-scheduler: The scheduling space grows exponentially with the increase in the number of input tensors in a kernel. In such cases, choosing an optimal schedule becomes a computationally intensive task. Protocolized Concrete Index Notation (CIN-P) (Ahrens et al., 2022), proposes an automated scheduler that enumerates every schedule of minimum depth and relies on the kernel being small. The cost analysis is done offline and programs with optimal schedules are run with uniform inputs to pick the best-performing one. CIN-P focuses solely on asymptotic costs and CIN-P for TACO discards schedules involving intermediate tensors of more than one dimension . SpTTN-Cyclops on the other hand tunes over both contraction path and index orderings for a given kernel. The pruning strategy restricts only the buffer size, not its dimensions. We also show that for analytic metrics of performance such as buffer size, more efficient search schemes are possible. Tensor Contraction Engine (TCE) (Baumgartner et al., 2005) automatically generates sequence of tensor contractions that minimize intermediate tensor sizes. It primarily focuses on dense tensor operations that are common in quantum chemistry computations.
@@
Inspector-executor models incorporated in the compiler transformation frameworks such as Sparse Polyhedral Framework (SPF) (Strout et al., 2016, 2018) enable optimization of sparse computations. For example, at runtime, an inspector can determine data access patterns and derive efficient schedules, which the executor can then invoke during computation. In (Zhao et al., 2022), the authors extend SPF to generate optimized sparse tensor codes and develop techniques for efficient co-iteration of sparse tensors. They focus on kernels that handle multiple sparse tensors and differ from the SpTTN kernels.
@@
General distributed-memory frameworks: DISTAL (Yadav et al., 2022a) extends TACO to target distributed systems. SpDISTAL (Yadav et al., 2022b) adopts single-node transformations of TACO and extends DISTAL’s distributed scheduling transformations with new constructs for describing distributions of sparse tensors. Each dimension of the tensor is partitioned and mapped to an n-dimensional abstract machine grid by matching the index names as specified by the user. SpDISTAL inherits the limitations of TACO in terms of finding an optimal code path for SpTTN kernels. Also, our framework provides automatic distributed memory parallelization without any user intervention. Deinsum (Ziogas et al., 2022) provides automatic distributed-memory parallelization of operations on tensors. Deinsum supports only dense tensors, and computes data movement-optimal schedules. TiledArray (Calvin et al., 2015; Calvin and Valeev, 2023) is a distributed-memory framework that natively supports block-sparse tensors.
Specialized library implementation for SpTTN kernels: SPLATT (Smith et al., 2015) provides an optimized implementation of MTTKRP on shared and distributed memory systems. GigaTensor (Kang et al., 2012) implements MTTKRP as a series of Hadamard products and uses the MapReduce paradigm. A parallel algorithm for TTMc which leverages multiple CSF representations is proposed in (Smith and Karypis, 2017). Parallel Tensor Infrastructure (ParTI!) (Li et al., 2018) is a library for sparse tensor operations (including MTTKRP) and tensor decompositions on multicore CPU and GPU architectures. In (Li et al., 2019), as part of ParTI!, the authors propose techniques to reorder the sparse tensor to improve the performance of MTTKRP. Methods to eliminate unnecessary communication when executing Sampled Dense-Dense Matrix Multiplication (SDDMM) and Sparse-Times-Dense Matrix Multiplication (SpMM) kernels in sequence is proposed in FusedMM (Bharadwaj et al., 2022).
@@ 7. Evaluation
All results are collected on the Stampede2 supercomputer. Each node has an Intel Xeon Phi 7250 CPU (“Knights Landing”) with 68 cores, 96GB of DDR4 RAM, and 16GB of high-speed on-chip MCDRAM memory. In our distributed memory experiments, we use 64 MPI processes per node. We select a loop nest with the maximum independent dense loops by imposing a bound on the intermediate tensor dimension, maintaining it at two. We compare SpTTN-Cyclops with TACO (Kjolstad et al., 2017) (commit ID 2b8ece4), general sparse pairwise contraction with CTF (Solomonik et al., 2014) (v1.5.5, commit ID 36b1f6d), SPLATT (Smith et al., 2015) (v1.1.0, commit ID 6cb8628) and SparseLNR (Dias et al., 2022) (branch dev-fuse, commit ID 8fafdd1). We present results for single node performance comparing with CTF, TACO, SparseLNR and SPLATT. For distributed memory performance we compare against CTF and SPLATT.
Figure 4. Single node performance of MTTKRP with $R=64$.
@@
Figure 6. (a) Strong scaling of TTTc on order 6 tensors of sparsity $1\%$ and $0.1\%$. $N$ and $R$ are set to $80$ and $16$, respectively. (b) Strong scaling of MTTKRP on nell-2 tensor with $R=64$. (c) Computation of TTMc using different index orders on order 3 tensor of sparsity $0.1\%$, $R=32$ and $N=8192$.
@@
Datasets: To evaluate the kernels, we use sparse tensors from Formidable Repository of Open Sparse Tensors and Tools (FROSTT) (Smith et al., 2017) and 1998 DARPA Intrusion Detection Evaluation (Jeon et al., 2015). For further analysis, we generate random sparse tensors with various dimensions and sparsities. The dense tensors are populated with random data / nonzero positions. If a tensor has identical dimensions, $N$ is used to represent size of the dimensions. A dense tensor shares some of its indices with the sparse tensor. The dimensions of the non-shared indices are denoted using $R$. For TTTc on order 6 tensors, we choose $R=16$.
@@
MTTKRP: In Figure 4 we compare the single node performance of SpTTN-Cyclops to that of the other frameworks. One of the optimal schedules to compute an order 3 MTTKRP in Equation 1 is to have a loop nest that partially contracts $\mathcal{T}$ with $\mathcal{U}$, and then with $\mathcal{V}$. This reduces the number of operations when compared to an unfactorized approach of TACO. SpTTN-Cyclops and SPLATT implement this factorize-and-fuse approach. We observe speedups of $1.3$x to $3.4$x when compared to TACO. SparseLNR fails to fuse loops within a kernel and has similar performance as TACO. We see speedups of $1.5$x to $1.7$x, and slowdowns of $0.8$x and $0.7$x when compared to SPLATT. These performance numbers can be attributed to code level optimizations, for example, compiler assisted loop generation (via metaprogramming) in SpTTN-Cyclops. CTF performs poorly when computing MTTKRP across all tensors. In Figure 5(b), we present strong scaling results for MTTKRP using the nell-2 tensor and compare it with SPLATT. Note that for this comparison, we account only for the compute times in both the frameworks. Despite being generic, our approach has performance close to a library implementation which is optimized for a specific kernel.
@@
TTMc: We observe substantial speedups over TACO and SparseLNR. Since we factorize the kernel into pairwise contractions and then fuse loops in SpTTN-Cyclops, there is an asymptotic reduction in computation complexity which translates to these observed speedups. We are able to run TTMc with TACO and SparseLNR only on two of the considered tensors, nell-2 and vast-3d. On nell-2, we observe a speedup of $29.3$x and $110.5$x over TACO and SparseLNR, respectively. Similarly, in vast-3d, we observe a speedup of $125.9$x and $4$x. We observe speedups over CTF in the range of $0.8$x to $12.6$x for the tensors considered (we are unable to run TTMc with CTF on enron and nell-2 tensors). On the nips tensor where the combination of the imbalanced dimensions of the tensor and the specific value of $R$ does not benefit the fused approach of SpTTN-Cyclops, we see a slowdown of $0.8$x. We are unable to execute TTMc on darpa using any of the approaches including SpTTN-Cyclops because of the larger memory footprint requirement for the contraction that cannot be accommodated on a single node.
In Figures 6(a) and 6(b), we present strong scaling results for MTTKRP and TTMc, respectively. SpTTN-Cyclops outperforms CTF on all node counts, and shows good scaling for both the kernels.
@@
TTTP: We present strong scaling results for TTTP in Figure 6(c). The single node performance of SpTTN-Cyclops over CTF is substantial with over $340$x speedup. We generate tensors with varied sparsity and fixed dimensions. We observe good scaling for all the considered tensors.
@@
TTTc: In Figure 5(a) we show strong scaling of the TTTc kernel. We generate two tensors of similar dimensions but with different sparsities. In both, SpTTN-Cyclops achieves good scaling. We are unable to run TTTc implementation in TACO and SparseLNR on these kernels. However, we generated a smaller tensor with dimensions $N=40$ and sparsity at $0.1\%$. SpTTN-Cyclops achieves a speedup of $534$x over TACO when computing TTTc on this tensor.
Impact of dense loops: Figure 5(c) shows the TTMc kernel performance for different index orderings. The contraction path for computing the kernel is $((\mathcal{T}\cdot\mathcal{V}\rightarrow\mathcal{X}),(\mathcal{X}\cdot\mathcal{U}\rightarrow\mathcal{S}))$. Index orders $(i,j,k,s)$ and $(i,j,s,r)$ for the two terms in the contraction path, respectively, require a scalar intermediate for the computation of the kernel. Instead, if we choose index order $(i,j,s,k)$ in the first term, we require a vector intermediate (See Section 3.3). Though both computation schedules scale, the latter index order achieves higher performance despite having a larger memory footprint. The contractions are offloaded to xAXPY (BLAS-1) (manually-implemented) and xGER (BLAS-2) kernels. In the index order that requires a scalar intermediate, only the contraction of $\mathcal{X}$ with $\mathcal{U}$ can be offloaded to a manually-implemented xAXPY kernel. Contraction of $\mathcal{T}$ with $\mathcal{V}$ is using an innermost sparse loop which affects performance.

### v2

@@ 5.1. Algorithm to Generate and Execute Loop Nests
1:Input: Sequence of terms that represent the contraction path. Each term is a set of three tensors, inp1, inp2 and op.
2:Input: Depth initially set to 0.
3:Output: Loop nest to compute the given kernel.
4:procedure LOOP_NEST(sequence_of_terms,
5:                              depth)
@@
8:         $\text{contract}(t.\text{inp1},t.\text{inp2},t.\text{op})$
@@
17:                       $b\leftarrow\text{buf\_terms}[i]$
@@ 5.2. Data Distribution
5.3. Example SpTTN Execution
In Figure 3, we show a fully fused loop nest for the order 4 TTMc kernel, $\mathcal{S}(i,r,s,t)=\sum_{j,k,l}\mathcal{T}(i,j,k,l)\cdot\mathcal{U}(j,r)\cdot\mathcal{V}(k,s)\cdot\mathcal{W}(l,t)$.
@@ 6. Related Work
SparseLNR (Dias et al., 2022) and ReACT (Zhou et al., 2023) extend TACO and COMET, respectively, with kernel distribution/fusion to support the factorize-and-fuse approach. The contraction path and loop orders for these loop nests are user-specified. Our main contribution is in fully enumerating the space of loop nests and finding a cost-optimal schedule automatically. Furthermore, in our evaluation (in Section 7), we show that SpTTN-Cyclops outperforms SparseLNR by orders of magnitude. For example, across various input tensors considered, SpTTN-Cyclops outperforms SparseLNR by 1.3x to 3.4x and 4x to 110.5x on MTTKRP and TTMc kernels, respectively.
Auto-scheduler: Tensor Contraction Engine (TCE) (Baumgartner et al., 2005) automatically generates sequence of tensor contractions that minimize intermediate tensor sizes. It primarily focuses on dense tensor operations that are common in quantum chemistry computations. The dynamic programming approach in TCE (Lam et al., 2000; Gao et al., 2007) adopts a bottom-up approach i.e., to find an optimal loop structure, the subtrees of the loop nest tree are evaluated first and memoized. Subsequently, at the root node, various loop structures including the possibility of fusing the subtrees are evaluated to pick the optimal loop structure. Furthermore, in TCE, the tree is partitioned into sub-problems by identifying a set of cut-points. There can be multiple cut-points at a given level. In SpTTN-Cyclops, at any given iteration, we split the problem into two sub-problems, i.e., only the first cut-point is considered, and the cost of the sub-problems is memoized. So a subproblem is a choice for the root index and prefixes of terms that contain that index to iterate over. This approach of SpTTN-Cyclops reduces the cost (for finding an optimal loop nest) when compared to choosing an index for each subtree at a given level and translates into better search complexity.
Protocolized Concrete Index Notation (CIN-P) (Ahrens et al., 2022), proposes an automated scheduler that enumerates every schedule of minimum depth and relies on the kernel being small. CIN-P focuses solely on asymptotic costs and CIN-P for TACO discards schedules involving intermediate tensors of more than one dimension. SpTTN-Cyclops on the other hand tunes over both contraction path and loop orderings. WACO (Won et al., 2023) co-optimizes the format and schedule of sparse tensor kernels using a sparse convolutional neural network to model and predict the runtime performance based on the sparsity patterns, formats, and schedules. SparseAuto (Dias et al., 2024) prunes the search space of schedules for sparse tensor contractions based on both time and intermediate tensor memory requirements. It uses Satisfiability Module Theory (SMT) solvers to pick the smallest number of possible schedules based on user-defined constraints. In CoNST (Raje et al., 2024), the authors use a constraint-based approach with a Z3 SMT solver to optimize schedules for sparse tensor contractions.
@@
Inspector-executor models incorporated in the compiler transformation frameworks such as Sparse Polyhedral Framework (SPF) (Strout et al., 2016, 2018) enable optimization of sparse computations. In (Zhao et al., 2022), the authors extend SPF to generate optimized sparse tensor codes. They focus on kernels that handle multiple sparse tensors and not SpTTN kernels.
@@
General distributed-memory frameworks: DISTAL (Yadav et al., 2022a) extends TACO to target distributed systems. SpDISTAL (Yadav et al., 2022b) adopts single-node transformations of TACO and extends DISTAL with new constructs for describing distributions of sparse tensors. SpDISTAL inherits the limitations of TACO in terms of finding an optimal code path for SpTTN kernels. Also, our framework provides automatic distributed memory parallelization without any user intervention. Deinsum (Ziogas et al., 2022) provides automatic distributed-memory parallelization of operations on dense tensors. TiledArray (Calvin et al., 2015; Calvin and Valeev, 2023) is a distributed-memory framework for block-sparse tensors.
[…]
@@ 7. Evaluation
All results are collected on the Stampede2 supercomputer. Each node has an Intel Xeon Phi 7250 CPU (“Knights Landing”) with 68 cores, 96GB of DDR4 RAM, and 16GB of high-speed on-chip MCDRAM memory. Additionally, we also run our experiments on a single node equipped with an Intel Xeon Silver 4314 processor (”Ice Lake”), which features a 64KB L1 data cache per core, 1MB L2 cache per core, and a 24MB shared L3 cache. The results of these experiments are reported in Figures 6 and 7. In our distributed memory experiments, we use 64 MPI processes per node. We select a loop nest with the maximum independent dense loops by imposing a bound on the intermediate tensor dimension, maintaining it at two. We compare SpTTN-Cyclops with TACO (Kjolstad et al., 2017) (commit ID 2b8ece4), general sparse pairwise contraction with CTF (Solomonik et al., 2014) (v1.5.5, commit ID 36b1f6d), SPLATT (Smith et al., 2015) (v1.1.0, commit ID 6cb8628) and SparseLNR (Dias et al., 2022) (branch dev-fuse, commit ID 8fafdd1). We present results for single thread performance comparing with CTF, TACO, SparseLNR and SPLATT. In SparseLNR, we try to use the optimal schedule derived from SpTTN-Cyclops, using its directives for kernel distribution and loop fusion. In TACO, we use the contraction path picked by SpTTN-Cyclops. For distributed memory performance we compare against CTF and SPLATT.
Figure 4. Single thread performance of MTTKRP with $R=64$.
@@
Datasets: To evaluate the kernels, we use sparse tensors from Formidable Repository of Open Sparse Tensors and Tools (FROSTT) (Smith et al., 2017) and 1998 DARPA Intrusion Detection Evaluation (Jeon et al., 2015). For further analysis, we generate random sparse tensors with various dimensions and sparsities. The dense tensors are populated with random data / nonzero positions. If a tensor has identical dimensions, $N$ is used to represent size of the dimensions. A dense tensor shares some of its indices with the sparse tensor. The dimensions of the non-shared indices are denoted using $R$.
@@
MTTKRP: In Figure 4, we compare the single thread performance of SpTTN-Cyclops to that of the other frameworks. One of the optimal schedules to compute an order 3 MTTKRP in Equation 1 is to have a loop nest that partially contracts $\mathcal{T}$ with $\mathcal{U}$, and then with $\mathcal{V}$. This reduces the number of operations when compared to an unfactorized approach of TACO. SpTTN-Cyclops and SPLATT implement this factorize-and-fuse approach. We observe speedups of $1.3$x to $3.4$x when compared to TACO. SparseLNR fails to fuse loops for this kernel and has similar performance to TACO. SpTTN-Cyclops achieves speedups of $1.5$x to $1.7$x, and slowdowns of $0.8$x and $0.7$x when compared to SPLATT. CTF performs poorly when computing MTTKRP across all tensors. We also conduct strong scaling experiments for MTTKRP on the nell-2 tensor. Despite being generic, our approach has performance close to SPLATT, a library optimized for a specific kernel.
TTMc: We observe substantial speedups over TACO. Since we factorize the kernel into pairwise contractions and then fuse loops in SpTTN-Cyclops, there is an asymptotic reduction in computation complexity which translates to these observed speedups. For an order 3 TTMc kernel (Equation 2), SparseLNR generates a schedule that contracts $\mathcal{T}$ with $\mathcal{U}$, and the result with $\mathcal{V}$. Only index $i$ is fused across the two pairwise contractions, and requires an intermediate tensor of $K\times R$ dimension. If the input tensor expression is $\mathcal{S}(i,r,s)=\sum_{j,k}\mathcal{T}(i,j,k)\cdot\mathcal{V}(k,s)\cdot\mathcal{U}(j,r)$, i.e., the position of $\mathcal{U}$ and $\mathcal{V}$ are interchanged in the expression, then SparseLNR defaults to the unfactorized approach of TACO. SpTTN-Cyclops generates a schedule that contracts $\mathcal{T}$ with $\mathcal{V}$, and the result with $\mathcal{U}$. Indices $i$ and $j$ are fused, and the intermediate tensor dimension is $S$ (Listing LABEL:lst:ttmc_fusion).
For an order 4 TTMc kernel (Section 5.3), SparseLNR generates a schedule that contracts tensors $\mathcal{T}$, $\mathcal{U}$ and $\mathcal{V}$ all-at-once, and the resulting intermediate tensor with $\mathcal{W}$. The intermediate tensor dimension is $L\times R\times S$. Only index $i$ is fused across these two contractions. The maximum loop depth is six. SpTTN-Cyclops generates an asymptotically optimal schedule as shown in Figure 3, which has a maximum loop depth of five.
@@
We are able to run TTMc with TACO and SparseLNR only on two of the considered tensors, nell-2 and vast-3d. On nell-2, we observe a speedup of $29.3$x and $110.5$x over TACO and SparseLNR, respectively. Similarly, in vast-3d, we observe a speedup of $125.9$x and $4$x. We observe speedups over CTF in the range of $0.8$x to $12.6$x for the tensors considered (we are unable to run TTMc with CTF on enron and nell-2 tensors). On the nips tensor where the combination of the imbalanced dimensions of the tensor and the specific value of $R$ does not benefit the fused approach of SpTTN-Cyclops, we see a slowdown of $0.8$x. We are unable to execute TTMc on darpa using any of the approaches including SpTTN-Cyclops because of the larger memory footprint requirement for the contraction that cannot be accommodated on a single node. In Figures 5(a) and 5(b), we present strong scaling results for MTTKRP and TTMc, respectively. SpTTN-Cyclops outperforms CTF on all node counts, and shows good scaling for both the kernels.
@@
TTTP: We present strong scaling results for TTTP in Figure 5(c). The single node performance of SpTTN-Cyclops over CTF is substantial with over $340$x speedup. We observe good scaling for all the considered tensors.
@@
TTTc: In our strong scaling analysis of TTTc, we evaluate two tensors of dimension $80$ ($R=16$) and sparsity at $1\%$ and $0.1\%$. In both, SpTTN-Cyclops achieves good scaling. SparseLNR generates a default TACO schedule for this kernel. We are unable to run TTTc implementation in TACO and SparseLNR on these kernels with the considered dimensions. However, we generated a smaller tensor with dimensions $N=40$ and sparsity at $0.1\%$. SpTTN-Cyclops achieves a speedup of $534$x over TACO on it.
Impact of intermediate tensor dimension: Consider an order 3 all-mode TTMc kernel, $\mathcal{S}(r,s,t)=\sum_{i,j,k}\mathcal{T}(i,j,k)\cdot\mathcal{U}(i,r)\cdot\mathcal{V}(j,s)\cdot\mathcal{W}(k,t)$ (all sparse indices are contracted). The contraction path chosen by SpTTN-Cyclops is $((\mathcal{T}_{ijk}\cdot\mathcal{W}_{kt}\rightarrow\mathcal{X}_{ijt}),(\mathcal{X}_{ijt}\cdot\mathcal{V}_{js}\rightarrow\mathcal{Y}_{ist}),(\mathcal{Y}_{ist}\cdot\mathcal{U}_{ir}\rightarrow\mathcal{S}_{rst}))$. For the chosen contraction path, if we consider a bound of two on the intermediate tensor dimension, the loop nest generated by SpTTN-Cyclops, $((i,j,k,t),(i,j,s,t),(i,r,s,t))$, has intermediate tensors $\mathcal{X}$ of size $T$ and $\mathcal{Y}$ of size $S\times T$. For the same contraction path, if we consider a bound of one on the intermediate tensor dimension, the loop nest generated, $((i,t,j,k),(i,t,j,s),(i,t,r,s))$, has intermediate tensors $\mathcal{X}$ of size $1$ (scalar) and $\mathcal{Y}$ of size $S$. In Figure 6, we show the single thread performance of the two loop nests generated by SpTTN-Cyclops for the order 3 all-mode TTMc kernel. We observe that the loop nest with intermediate tensors of size $T$ and $S\times T$ performs better than the loop nest with intermediate tensors of size $1$ and $S$, despite having a larger memory footprint. The contractions in Loop Nest #2 are offloaded to two xAXPY (BLAS-1) (manually-implemented) and one xGER (BLAS-2) kernels. Loop Nest #1, on the other hand, employs an innermost sparse loop to compute the intermediate tensor $\mathcal{X}$. Consequently, only two BLAS kernels are used in this loop nest: one for computing $\mathcal{Y}$ and the other for $\mathcal{S}$.
Impact of loop order: For the order 3 all-mode TTMc kernel and the contraction path chosen by SpTTN-Cyclops, we randomly select $25\%$ of all possible loop orders that have the sparse indices iterated over in the order in which they are stored in the CSF tree. In Figure 7, we show the single thread performance of these randomly picked loop orders. In the loop order picked by SpTTN-Cyclops the intermediate tensors are within a considerable memory bound and also allows for the maximum use of BLAS kernels.
Figure 6. Single thread performance of an order 3 all-mode TTMc contraction. Loop Nest #1 has a bound of 1 and Loop Nest #2 has a bound of 2 on the intermediate tensor dimension. $R=64$.
Figure 7. Single thread performance of an order 3 all-mode TTMc contraction with $N=1024$, $R=32$, and sparsity at $0.1\%$ using randomly picked loop orders. The red line represents the cut-off and the green line represents the runtime of the loop order picked by SpTTN-Cyclops.